A 35-L block of ice (density at 0.9167 g/cm3) sits on a surface that is 25 inches wide and 16 inches long on Mars, where the acceleration due to gravity is only 38.74% that of Earth’s. Express the pressure exerted on this surface using atmospheres.

1
Expert's answer
2021-03-18T05:57:17-0400

16 in = 0.4064 m.

25 in = 0.635 m.

p= m/V. m = p x V.

P= F/S.

F= mg.

S= a x b.

m=35 x 0.9167 = 32,0845 kg.

F= 32,0845 x 10 x 0.3874 = 124,295353 N.

S= 0.635 x 0.4064 = 0,258064 m2

P= 124,295353/0,258064 = 481,645456 Pascal.